### v1

@@ III-C Ablation Study
In addition to the ablation study in Table III, we further conduct an ablation study to analyze the effectiveness of our method on the entire validation set. As shown in IV, after adding the proposed method, an improvements can be seen on both the different occlusion subset and the entire validation set, especially $\text{PQ}^{\text{St}}$ improves by 0.8 on the entire validatoin set. This indicates that proposed contrastive learning method enables the model to learn a more occlusion-robust feature, thereby improving its overall performance.
Then, we investigate the influence of the margins in our proposed contrastive learning in Table V. Particularly, we first evaluate the threshold combination of $\tau_{l,h}$ and $\tau_{m}$ by keeping one threshold fixed while increasing the other. When the first threshold $\tau_{l,h}$ is fixed at 0.3, increasing the second threshold $\tau_{m}$ (from 0.4 to 0.8) shows minimal effect on overall performance. When $\tau_{l,h}$ is increased to 0.4, the combination of $\tau_{l,h}=0.4$ and $\tau_{m}=0.6$ achieves the highest PQ. However, under further increases of $\tau_{m}$ beyond 0.6, the overall performance starts to decline. This trend demonstrates that neglecting the influence of either threshold will result in a degradation of overall performance.
Our method is the initial attempt to utilize the occlusion annotations. We believe that a more significant improvement can be achieved when further advanced methods are proposed to take better advantange of our proposed dataset.
@@ IV Conclusion
In this work, we introduce COCO-Occ, a novel large-scale dataset designed to facilitate research on occlusion in panoptic segmentation and other image understanding tasks. The dataset, derived from the COCO dataset with manual occlusion annotations, provides a comprehensive benchmark with over 30,000 training and 5,000 test images categorized into three occlusion levels. Our experiments systematically evaluated the performance of SOTA panoptic segmentation methods on this dataset, revealing that occlusion, especially at higher levels, substantially degrades model performance. Additionally, we proposed a contrastive learning-based method to enhance model robustness in handling occlusion. Despite its simplicity, our approach yielded performance improvements over the baseline. This work not only highlights the significant impact of occlusion on panoptic segmentation task but offers a quantitative foundation for future research into more effective occlusion-aware methods.

### v2

@@ III-C Ablation Study
In addition to the ablation study in Table III, we further conduct an ablation study to analyse the effectiveness of our method on the entire validation set.
As shown in IV, after adding the proposed method, an improvement can be seen on both the various occlusion subsets and the entire validation set, especially $\text{PQ}^{\text{St}}$, which improves by 0.8 on the entire validation set. This indicates that the proposed contrastive learning method enhances the model’s ability to learn a more occlusion-robust feature, thereby improving its overall performance.
Then, we investigate the influence of the margins in our proposed contrastive learning method in Table V. In particular, we first evaluate the combination of thresholds of $\tau_{l,h}$ and $\tau_{m}$ by keeping one threshold fixed while increasing the other. When the first threshold $\tau_{l,h}$ is fixed at 0.3, increasing the second threshold $\tau_{m}$ (from 0.4 to 0.8) shows a minimal effect on overall performance. When $\tau_{l,h}$ increases to 0.4, the combination of $\tau_{l,h}=0.4$ and $\tau_{m}=0.6$ achieves the highest PQ. However, with further increases of $\tau_{m}$ beyond 0.6, the overall performance starts to decline. This trend demonstrates that neglecting the influence of either threshold will result in a degradation of overall performance.
Our method is the initial attempt to utilise the occlusion annotations. We believe that a more significant improvement can be achieved when further advanced methods are proposed to better optimise our proposed dataset.
@@ IV Conclusion
In this work, we introduce COCO-OLAC, a novel large-scale dataset designed to facilitate research on occlusion in panoptic segmentation and other image understanding tasks. The dataset, derived from the COCO dataset with manual occlusion annotations, provides a comprehensive benchmark with over 30,000 training images and 5,000 test images categorised into three occlusion levels. Our experiments systematically evaluated the performance of SOTA panoptic segmentation methods on this dataset, revealing that occlusion, especially at higher levels, substantially degrades model performance. In addition, we propose a contrastive learning-based method to enhance model robustness in handling occlusion. Despite its simplicity, our approach yields performance improvements over the baseline. This work not only highlights the significant impact of occlusion on the panoptic segmentation task but also offers a quantitative foundation for future research into more effective occlusion-aware methods.